### v2

# Thought Anchors: Which LLM Reasoning Steps Matter?

Abstract
Reasoning large language models have recently achieved state-of-the-art performance in many fields. However, their long-form chain-of-thought reasoning creates interpretability challenges as each generated token depends on all previous ones, making the computation harder to decompose. We argue that analyzing reasoning traces at the sentence level is a promising approach to understanding reasoning processes. We present three complementary attribution methods: (1) a black-box method measuring each sentence’s counterfactual importance by comparing final answers across 100 rollouts conditioned on the model generating that sentence or one with a different meaning; (2) a white-box method of aggregating attention patterns between pairs of sentences, which identified “broadcasting” sentences that receive disproportionate attention from all future sentences via “receiver” attention heads; (3) a causal attribution method measuring logical connections between sentences by suppressing attention toward one sentence and measuring the effect on each future sentence’s tokens. Each method provides evidence for the existence of thought anchors, reasoning steps that have outsized importance and that disproportionately influence the subsequent reasoning process. These thought anchors are typically planning or backtracking sentences. We provide an open-source tool (thought-anchors.com) for visualizing the outputs of our methods, and present a case study showing converging patterns across methods that map how a model performs multi-step reasoning. The consistency across methods demonstrates the potential of sentence-level analysis for a deeper understanding of reasoning models.
1 Introduction
Figure 1: Summary of our three methods for principled attribution to important sentences in reasoning traces. A. An example reasoning trace with sentences labeled per our taxonomy. B. Our proposed methods are: black-box resampling, receiver heads, and attention suppression. C. A directed acyclic graph among sentences prepared by one of our techniques, made available open source.
Training large language models to reason with chain-of-thought [28, 23, 34] has led to significant advances in capabilities [26]. The resulting reasoning traces are already used in safety research [3, 29]. However, there has been little work adapting interpretability methods to this new paradigm (though see [31, 14]). Traditional mechanistic interpretability [25, 24] methods often focus on a single forward pass of the model: understanding the activations, how they are processed by each layer, and how they are converted to the final output [33, 16, 15]. However, this framework is too fine-grained for autoregressive reasoning models, which consume their own output tokens.
A core step in many interpretability strategies is to decompose the model into smaller parts that can be analyzed independently [20]. A natural decomposition for chain-of-thought is into individual sentences and how they depend on each other. Interpretations of neural network behavior operate at varying levels of abstraction [11, 12], and sentence-level explanations strike an intermediate abstraction depth. Compared to tokens, sentences are more coherent and often coincide with reasoning steps extracted by an LLM [31, 2]. Compared to paragraphs, sentences are less likely to conflate reasoning steps and may serve as an effective target for linking different steps.
Prior work has established that different sentences within reasoning traces perform distinct functions. Backtracking sentences (e.g., “Wait…”) cause the model to revisit earlier conclusions, which boosts final-answer accuracy [22]. Other research has distinguished sentences based on whether they retrieve new information or execute deduction with existing information [31]. Hence, reasoning may follow an overarching structure, where sentences can introduce and pursue high-level computational goals. Yet, approaches for mapping this structure remain limited.
We propose three complementary methods for mapping the structure of reasoning traces that focus on what we term thought anchors: critical reasoning steps that guide the rest of the reasoning trace. We provide evidence for this type of anchoring based on black-box evidence from resampling and white-box evidence based on attention patterns.
First, in section 3 we present a black-box method for measuring the counterfactual impact of a sentence on the model’s final answer and future sentences. We repeatedly resample reasoning traces from the start of each sentence. Based on resampling data, we can quantify the impact of each sentence on the likelihood of any final answer or the likelihood of producing any subsequent sentence. Via this resampling approach, we can additionally distinguish planning sentences that initiate computations leading to some answer from sentences performing computations necessary for the answer but which are predetermined.
Second, in section 4 we present a white-box method for evaluating importance based on the sentences most attended. Our analyses reveal “receiver” heads that narrow attention toward particular past “broadcasting” sentences. Compared to base models, where attention is more diffuse, reasoning models display overall greater attentional narrowing through receiver heads, and these heads have an outsized impact on the model’s final answer. We develop a systematic approach to identifying receiver heads and show how evaluating sentences on the extent to which they are broadcast by these heads provides a mechanistic measure of importance.
Third and finally, in section 5 we present a method that measures the causal dependency between specific pairs of sentences in a reasoning trace. For each sentence in a trace, we intervene by masking all attention to it from subsequent tokens. We then measure the effect of suppression on subsequent token logits (KL divergence) compared to those generated during the absence of suppression. Averaging token effects by sentence, this strategy measures each sentence’s direct causal effect on each subsequent sentence.
Applying these techniques, our work suggests that analyzing reasoning through sentence-level units introduces new domains through which reasoning models can be understood. Our work also opens the door to more precise debugging of reasoning failures, identification of sources of unreliability, and the development of techniques to enhance the reliability of reasoning models.
2 Setup
2.1 Sentence taxonomy for reasoning functions
We categorized sentences using the framework by [31], which describes distinct reasoning functions within a reasoning trace. We define eight categories (see examples and frequencies in Appendix A):
  1. 1.
    Problem Setup: Parsing or rephrasing the problem
  2. 2.
    Plan Generation: Stating or deciding on a plan of action, meta-reasoning
  3. 3.
    Fact Retrieval: Recalling facts, formulas, problem details without computation
  4. 4.
    Active Computation: Algebra, calculations, or other manipulations toward the answer
  5. 5.
    Uncertainty Management: Expressing confusion, re-evaluating, including backtracking
  6. 6.
    Result Consolidation: Aggregating intermediate results, summarizing, or preparing
  7. 7.
    Self Checking: Verifying previous steps, checking calculations, and re-confirmations
  8. 8.
    Final Answer Emission: Explicitly stating the final answer
Each sentence in the analyzed response is assigned to one of these categories using an LLM-based auto-labeling approach. The prompt used to label sentence categories is provided in Appendix B. Some categories that rarely appear are omitted from the figures below. Probing classifiers achieve high accuracy in distinguishing sentences and their corresponding categories (see Appendix C).
2.2 Model and dataset
Our analyses are based on the DeepSeek R1-Distill Qwen-14B model (48 layers) [8]. We used a temperature of $0.6$ and a top-p value of $0.95$ in all of our experiments. We used the MATH dataset [17] specifically targeting 10 challenging mathematics questions that the model correctly solves 25-75% of the time, identified by testing the model on 1,000 problems 10 times each. For each selected problem, we generated one correct and one incorrect reasoning trace, producing 20 responses. The average response is 144.2 sentences long (95% CI: [116.7, 171.8] and 4208 tokens (95% CI: [3479, 4937]). We focus only on sentences before the resampling method indicates the model has converged on an answer (i.e., after which it gives the same response in >98% of resamples). In Appendix D, we provide results from applying our techniques to other reasoning models.
3 Measuring counterfactual influence
Some sentences matter more than others, but which ones matter most depends on how we define and measure importance. We frame sentence-level importance as a question of counterfactual influence: how does including or excluding a sentence affect subsequent steps and the model’s final output?
3.1 Forced answer importance
[…]
The taxonomic categories reveal that active computation statements produce the greatest distributional shifts in this forced answer importance (Figure 3A). This aligns with prior findings that reasoning models update their beliefs throughout the trace, and suggests intermediate calculations alter the model’s answer responses. However, active computation steps may not provide a comprehensive view on importance, as they may be predetermined by upstream statements.
3.2 Counterfactual importance
A natural limitation of the forced-answer approach is that a sentence $S$ may be necessary for some final answer, but consistently produced by the LLM late in the reasoning trace. This means that forced answer accuracy will be low for all sentences that occur before $S$, not allowing the importance of these earlier steps to be accurately determined.
Consider a rollout which consists of sentences $S_{1},S_{2},\dots,S_{i},\dots,S_{M}$ and a final answer $A$. We use resampling to define a metric capturing the extent to which sentence $S$ causes $A$ to be incorrect. [^1] All $S_{j}$ and $T_{k}$ as well as $A$ in this section are random variables, since we sample with non-zero temperature. We call this metric counterfactual importance. We can motivate and define this metric with three steps:
  1. 1.
    Rollout sampling. For a given sentence $S_{i}$, we generate a distribution over final answers by generating 100 rollouts both without sentence $S_{i}$ (the base condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},T_{i},\dots,T_{N},A^{\prime}_{S_{i}}$), and another distribution with sentence $S$ (the intervention condition, with rollouts of the form $S_{1},S_{2},\dots,S_{i-1},S_{i},\dots,S_{M},A_{S_{i}}$).
  2. 2.
    Naive distributional comparison. We can compute the KL Divergence between the final answer distributions in the two conditions, i.e., $\text{importance}_{\text{r}}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}})||p(A_{S_{i}})]$, providing a scalar measure of how much sentence $S_{i}$ changes the answer. We call this the resampling importance metric.
  3. 3.
    Semantic filtering. The problem with resampling importance is that in rollout resampling, if $T_{i}$ is identical or similar to $S_{i}$ then we do not get much information about whether $S_{i}$ is important or not. Therefore, we write $S\not\approx T$ if two sentences $S$ and $T$ are dissimilar, defined as having cosine similarity less than the median value across all sentence pairs in our dataset, when fed into an embeddings model (see Appendix E for details). Therefore we can finally define the counterfactual importance by conditioning on $T_{i}\not\approx S_{i}$ i.e., $\text{importance}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}}|T_{i}\not\approx S_{i})||p(A_{S_{i}})]$.
Figure 2: Accuracy over 100 rollouts at each sentence for (A) one correct and (B) one incorrect base solution. Red dots mark significant spikes or dips. Local minima and maxima sentences are annotated with category initials (e.g., PG = plan generation). The analyses below focus on the counterfactual KL-divergence between sentences, but resampling accuracy is visualized here as it is more intuitive.
Since we resample all steps after a given sentence $S_{i}$, we avoid the aforementioned limitation of forced-answering. We also provide empirical evidence that the principled counterfactual importance definition in 1-3 above is useful, by comparing it to the resampling importance in Appendix G.
Across our dataset, we find that plan generation and uncertainty management (e.g., backtracking) sentences consistently show higher counterfactual importance than other categories like fact retrieval or active computation (see Figure 3B). This supports the view that high-level organizational sentences anchor, organize, and steer the reasoning trajectory. We argue that this approach provides more informative results compared to forced answer importance and prior token- or attention-based metrics.
Figure 3: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown (see Appendix F).
3.3 Sentence-to-sentence importance
In addition to estimating a sentence’s effect on the final answer, we extend our framework to quantify its causal influence on each individual future sentence. This sentence-level attribution helps uncover the structure of the model’s chain-of-thought: which sentences initiate subroutines, influence future claims, or act as attractors for subsequent steps. We estimate the counterfactual importance of sentence $S_{i}$ on a future sentence $S^{\text{Fut.}}$ formally with:
$\text{importance}(S_{i}\to S^{\text{Fut.}})=\mathbb{P}(S^{\text{Fut.}}\in_{\approx}\{S_{i},\dots,S_{M}\})-\mathbb{P}(S^{\text{Fut.}}\in_{\approx}\{T_{i},\dots,T_{N}\}|T_{i}\not\approx S_{i})$ (1)
Intuitively, on the right-hand side of Equation 1, the first term is the probability that a future sentence $S^{\text{Fut.}}$ will semantically occur given that $S_{i}$ was present in the trace, and the second term is the corresponding probability when $S_{i}$ is resampled with a non-equivalent sentence. A positive score indicates that sentence $S_{i}$ increases the likelihood of producing $S^{\text{Fut.}}$ (i.e., $S_{i}$ upregulates $S^{\text{Fut.}}$), while a negative score suggests that it suppresses or inhibits it.
In this context $S^{\text{Fut.}}$ semantically occurs if, when we extract the sentences and identify the best candidate match for $S^{\text{Fut.}}$ using cosine similarity between sentence embeddings, it has greater than $0.8$ cosine similarity (i.e., the median value in our dataset) to that sentence. Pseudocode for estimating sentence-to-sentence importance and empirical values of this metric can be found in Appendix G.
4 Attention aggregation
We hypothesize that important sentences may receive heightened attention from downstream sentences. Although attention weights do not necessarily imply causal links (see Section 5), heightened attention is a plausible mechanism by which important sentence may exert their influence on subsequent sentences. We conjecture further that a high focus on important sentences may be driven by specific attention heads, and by tracking such heads, we may pinpoint key sentences.
We assessed the degree different heads narrow attention toward particular sentences. First, for each reasoning trace, we averaged each attention head’s token-token attention weight matrix to form a sentence-sentence matrix, where each element is the mean across all pairs of tokens between two sentences. Based on each attention matrix, we computed the mean of its column below the diagonal to measure the extent each sentence receives attention from all downstream sentences; averaged only among pairs at least four sentences apart to focus on distant connections. This generates a distribution for each head (e.g., Figure 4A), and the extent each head generally narrows attention toward specific sentences can be quantified as its distribution’s kurtosis (computed for each reasoning trace, then averaged across traces). Plotting each head’s kurtosis reveals that some attention heads strongly narrow attention toward specific, possibly important, sentences in the reasoning trace (Figure 4B).
4.1 The identification of receiver heads
We refer to attention heads that narrow attention toward specific sentences as “receiver heads”. These heads are more common in later layers (Appendix H). To formally assess the existence of receiver heads, we tested whether some attention heads consistently operate in this role. Specifically, we measured the split-half reliability of heads’ kurtosis scores. We found a strong head-by-head correlation ($r$ = .67) between kurtosis scores computed for half of the problems with kurtosis scores for the other half of problems. Thus, some attention heads consistently operate as receiver heads, albeit with some heterogeneity across responses in which heads narrow attention most.
Figure 4: A. Lines show the vertical attention scores for each sentence by the 40 different heads in layer 36. Head 6 has been highlighted as a receiver head, and its corresponding attention weight matrix is shown for reference. Its prominent spikes cause the distribution to have a high kurtosis. B. Histogram of these kurtosis values across all attention heads, averaged across all reasoning traces.
Receiver heads usually direct attention toward the same sentences. Among the 16 heads with the highest kurtoses, we computed the sentence-by-sentence correlation between the vertical-attention scores for each pair of heads; calculated separately for each reasoning trace, then averaged. This produced an large correlation (mean $r$ = .60). Thus, receiver heads generally attend the same sentences, albeit with some heterogeneity (this correlation among randomly selected heads is $r$ = .45). This convergence is consistent with the existence of sentence importance, which these heads identify.
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models that enhances their performance. Comparing R1-Distill-Qwen-14B (reasoning) and Qwen-14B (base) suggests that the reasoning model’s receiver heads will narrow attention toward singular sentences to a greater degree (appendix I). Furthermore, ablating receiver heads leads to a greater reduction in accuracy than ablating self-attention heads at random (appendix J). Altogether, these findings are consistent with receiver heads and thought-anchor sentences playing special roles in reasoning.
4.2 Links to resampling importance and sentence types
[…]
Figure 5: The boxplot shows the average top-16 receiver-head score for each sentence type. The boxes correspond to the interquartile range across different reasoning traces.
5 Attention suppression
5.1 Approach
A natural concern with examining attention weights is that they do not measure causality. Further, receiver heads are not necessarily suited for identifying sentences’ links to individual future sentences. Hence, we next focused on causal attribution for sentence-sentence dependencies. The sentence-sentence analyses based on resampling also examined such dependencies but had limited precision for mapping logical connections, as counterfactual importance represents the total effect (direct and indirect) of one sentence on another. The present method instead aims to isolate the direct effects between sentences, which we expect to permit more precise modeling of logical connections.
We suppress all attention (all layers and heads) toward a given sentence and examine the impact on future sentences. The impact is defined as the KL divergence between a token’s logits and its baseline logits without suppression. The effect on a future sentence is calculated as the average of its constituent token log-KL divergences. This technique assumes (i) token logits accurately capture a sentence’s semantic content and (ii) suppressing attention does not problematically induce out-of-distribution behavior. Hence, we must validate correspondence between the attention-suppression measure and resampling measure, which, although less precise, does not hinge on these assumptions.
5.2 Correlations with the resampling-based importance matrix
[…]
6 Case study
The three presented techniques cover different aspects of attribution within a reasoning trace. We will now illustrate the utility and the complementary nature of these techniques by applying them to the model’s response for one problem: “When the base-16 number $66666_{16}$ is written in base 2, how many base-2 digits (bits) does it have?” (Figure 6) (see Appendix L for the full transcript).
Figure 6: Case study: problem #4682 (correct). Red matrix shows the effect of suppressing one sentence (x-axis) on a future sentence (y-axis). Darker colors indicate higher values. Bottom-left line plot shows the average attention toward each sentence by all subsequent sentences via the top-32 receiver heads (32 attention heads with the highest kurtosis score). Flowchart summarizes the model’s CoT with chunks defined around key sentences receiving high attention via receiver heads. Sentence 13 is emphasized as it has high counterfactual importance per the resampling method (see Figure 2A).
6.1 Resampling
The model approaches this problem by first considering that $66666_{16}$ contains five base-16 digits and any base-16 digit can be represented with four base-2 digits. Following this logic, the CoT starts by pursuing the answer “20 bits”. However, this initial answer overlooks that $6_{16}$ is $110_{2}$ (i.e., not $0110_{2}$ due to the leading zero), so the correct answer is actually “19 bits”. In sentence 13, the model pivots toward this correct answer by initiating computations to convert $66666_{16}$ to decimal and then binary.
The resampling method (section 3) reveals this initial poor trajectory and pivot upwards (shown previously in Figure 2A). Specifically, from sentences 6-12, expected accuracy steadily declines, but sentence 13 causes counterfactual accuracy to drastically increase. The pivotal role of 13 is notably missed if one instead evaluates accuracy by forcing the model to give a response immediately, as some prior studies have done, which instead yields 0
6.2 Receiver heads
The trajectory toward the final correct answer can be understood as a series of computational chunks (see flowchart in Figure 6). First, the model prepares a formula for converting $66666_{16}$ to decimal (sentences 13-19). Next, the model computes the answer to that formula, finding that $66666_{16}$ is 419,430 in decimal (sentences 20-33). The model subsequently converts that number to binary by putting forth another formula and solving it, $floor(log_{2}(419,430))+1=19$, to derive that the answer is “19 bits” (sentences 34-41). The model then notes a discrepancy with the earlier 20-bit solution (sentences 42-45). The model hence initiates new computations that verify that it computed the decimal value of $66666_{16}$ correctly (sentences 46-58) and that it computed the binary conversion accurately (sentences 59-62). Equipped with this increased certainty about 19-bit answer, the model discovers why its initial 20-bit idea was incorrect: “because leading zeros are not counted” (Sentence 66). This overall narrative is based on our analysis of attention patterns (section 4): Receiver attention heads pinpoint sentences initiating computations or stating key conclusions, thereby segmenting the reasoning trace into seemingly meaningful chunks (Figure 6).
6.3 Attention suppression
Along with being organized into computational chunks, the reasoning displays a scaffold related to sentence-sentence dependencies (Figure 6). One piece of this structure is a self-correction pattern involving an incorrect proposal, a detected discrepancy, and a final resolution. Specifically, the model initially proposes an incorrect answer of “20 bits”, which it decides to recheck (sentence 12). This leads to a discrepancy with the “19 bits” answer computed via decimal conversion (sentences 43 & 44). After rechecking its arithmetic supporting the “19 bit” answer, the model returns to the discrepancy (sentence 65) and then produces an explanation for why the “20 bits” answer is incorrect (sentence 66). This can be seen as a tentative CoT circuit, where two conclusions conflict to produce a discrepancy, which in turn encourages the model to resolve the discrepancy. Within this wide-spanning scaffold, there exist further dependencies, corresponding to verifying an earlier computation. Specifically, the model finishes computing the decimal value of $66666_{16}$ as 419,430 (sentence 32), later decides to verify that decimal conversion (sentence 46), and finally confirms that the original value is correct (sentence 59). This can be seen as further indication of CoT circuitry.
We identified these linkages based on the attention-suppression matrix (section 5), which contains local maxima at these linkages (12 → 43, 43 → 65, 12 → 66; 32 → 46, 32 → 59). Notice that many of the sentences pinpointed by the attention-suppression technique overlap with the sentences receiving high attention from receiver heads. Adding to the receiver-head conclusions, the attention suppression technique shows how information flows between these key sentences that structure the reasoning trace.
6.4 Open source interface
We released an open source interface (thought-anchors.com) for analyzing and visualizing reasoning traces. The reasoning trace can be visualized as an annotated directed acyclic graph using the importance metrics from Section 3, which may benefit interpretability and reasoning model debugging. We show our proof-of-concept interface in Figure 1C, where important sentences are represented by larger nodes and sentences that causally affect each other are connected with dashed gray lines. Hovering over a node lets users inspect sentence properties or compare alternative rollouts.
7 Related work
Reasoning advances in LLMs.
Chain-of-thought reasoning [34, 23, 28] has driven significant capabilities improvements in large language models. State-of-the-art performance on complex tasks comes from “thinking models” like DeepSeek-R1 [8], OpenAI’s o1 [26], and Claude 3.7 [1] are explicitly optimized for this with RLVR[36]. These increased reasoning capabilities introduce novel safety challenges, including potential unfaithfulness in the reasoning itself [19]. While current alignment evaluation methods often rely on examining reasoning traces [3, 29], we cannot rely on the model’s reasoning being faithful to the internal reasoning[7, 30, 2], creating a need for more mechanistic analyses.
Importance of individual steps. Many CoT sentences can be deleted or rewritten with limited impact on task accuracy [32, 21]. Prior attribution methods suggest that a subset of steps disproportionately drive the final answer: Shapley values [10], ROSCOE metrics [13], gradient-based scores [35], and resampling at fork tokens [4]. Complementing these, our approach provides a more principled framework for understanding how reasoning traces are constructed around key sentences.
White-box analyses of reasoning. Mechanistic studies locate specialised components that propagate information across steps: iteration or receiver heads [6], small sets of heads necessary for CoT performance [37], steerable activation directions for planning and backtracking [31], and deduction circuits in symbolic tasks [5]. Other work identifies phase shifts separating pre-training priors from contextual reasoning [9]. These findings motivate our focus on sentence-level structure rather than token-level saliency. However, they do not provide general methods for understanding the causal interdependence between sentences.
8 Discussion and Limitations
This work presents initial steps towards a principled decomposition of reasoning traces with a focus on identifying thought anchors: sentences with outsized importance on the model’s final response, specific future sentences, and downstream reasoning trajectory. We have also begun unpacking the attentional mechanisms associated with these important sentences. We expect that understanding thought anchors will be critical for interpreting reasoning models and ensuring their safety.
We view this as preliminary work. Our analyses require refinement to fully grapple with how downstream sentences may be overdetermined by different trajectories in a reasoning trace or independent sufficient causes. Additionally, we do not formally examine the role of error correction. Our receiver-head analyses are confounded by a sentence’s position in the reasoning trace (see Appendix M). Further, our attention-suppression work is limited because it effectively requires the model to process out-of-distribution information.
Nonetheless, we believe that we have demonstrated that our metrics are an advance on prior work, interrupting models and forcing final answers. The surprising degree of shared structure we have found across our three methods illustrates the potential value of future work in this area and points to the possibility of more powerful interpretability techniques to come.
Acknowledgments
This work was conducted as part of the ML Alignment & Theory Scholars (MATS) Program. We would like to thank Iván Arcuschin, Constantin Venhoff, and Samuel Marks for helpful discussions and feedback. We particularly thank Stefan Heimersheim for his valuable feedback and insightful suggestions, including ideas for experimental approaches that helped strengthen our analysis and contributed to the clarity of our presentation. We also thank members of Neel Nanda’s MATS stream for engaging brainstorming sessions, thoughtful questions during our presentations, and ongoing discussions that helped shape our approach.
Author Contributions
Both first authors, Paul C. Bogdan and Uzay Macar, contributed to the research, engineering, and writing of the paper. Neel Nanda was the main supervisor and provided feedback and guidance throughout the project. Arthur Conmy proposed the black-box resampling method for measuring the importance of sentences and also provided feedback and guidance throughout the project.
Appendix A Sentence taxonomy
Building on top of the framework presented by [31], we developed a taxonomy consisting of eight distinct sentence categories that capture reasoning functions in mathematical problem-solving. Each category represents a specific cognitive operation. The functions and examples for each category are given in Table 1. Notably, the uncertainty management category includes backtracking sentences.
Table 1: Sentence taxonomy with reasoning functions in problem-solving
Category Function Examples
Problem Setup Parsing or rephrasing the problem (e.g., initial reading) I need to find the area of a circle with radius 5 cm.
Plan Generation Stating or deciding on a plan of action, meta-reasoning I’ll solve this by applying the area formula.
Fact Retrieval Recalling facts, formulas, problem details without computation The formula for the area of a circle is $A=\pi r^{2}$.
Active Computation Algebra, calculations, or other manipulations toward the answer Substituting $r=5$: $A=\pi\times 5^{2}=25\pi$.
Uncertainty Management Expressing confusion, re-evaluating, including backtracking Wait, I made a mistake earlier. Let me reconsider…
Result Consolidation Aggregating intermediate results, summarizing, or preparing So the area is $25\pi$ square cm which is approximately…
Self Checking Verifying previous steps, checking calculations, and re-confirmations Let me verify: $\pi r^{2}=\pi\times 5^{2}=25\pi$. Correct.
Final Answer Emission Explicitly stating the final answer Therefore, the answer is…
The distribution of categories across our dataset as shown in Figure 7 reveals that active computation constitutes the largest proportion (32.7%), followed by fact retrieval (20.1%), plan generation (15.5%), and uncertainty management (14.0%). The sequential structure of reasoning is reflected in the rarity and positioning of problem setup (2.4%), which typically occurs at the beginning, and final answer emission (0.7%), which predominantly appears toward the end of the reasoning process.
Figure 7: Counts and frequencies of taxonomic sentence categories in our dataset.
Appendix B Prompt information
We used the following prompt with OpenAI GPT-4o (April-May, 2025) to annotate each sentence:
You are an expert in interpreting how LLMs solve math problems using multi-step reasoning. Your task is to analyze a chain-of-thought reasoning trace, broken into discrete text sentences, and label each sentence with:
1. **function_tags**: One or more labels that describe what this sentence is *doing* functionally in the reasoning process.
2. **depends_on**: A list of earlier sentence indices that this sentence directly depends on, e.g., uses information, results, or logic introduced in earlier sentences.
This annotation will be used to build a dependency graph and perform causal analysis, so please be precise and conservative: only mark a sentence as dependent on another if its reasoning clearly uses a previous sentence’s result or idea.
Function Tags:
1. problem_setup: Parsing or rephrasing the problem (initial reading or comprehension).
2. plan_generation: Stating or deciding on a plan of action (often meta-reasoning).
3. fact_retrieval: Recalling facts, formulas, problem details (without immediate computation).
4. active_computation: Performing algebra, calculations, manipulations toward the answer.
5. result_consolidation: Aggregating intermediate results, summarizing, or preparing final answer.
6. uncertainty_management: Expressing confusion, re-evaluating, proposing alternative plans (includes backtracking).
7. final_answer_emission: Explicit statement of the final boxed answer or earlier sentences that contain the final answer.
8. self_checking: Verifying previous steps, checking calculations, and re-confirmations.
9. unknown: Use only if the sentence does not fit any of the above tags or is purely stylistic or semantic.
Dependencies:
For each sentence, include a list of earlier sentence indices that the reasoning in this sentence *uses*. For example:
- If sentence 9 performs a computation based on a plan in sentence 4 and a recalled rule in sentence 5, then depends_on: [4, 5]
- If sentence 24 plugs in a final answer to verify correctness from sentence 23, then depends_on: [23]
- If there’s no clear dependency use an empty list: []
- If sentence 13 performs a computation based on information in sentence 11, which in turn uses information from sentence 7, then depends_on: [11, 7]
Important Notes:
- Make sure to include all dependencies for each sentence.
- Include both long-range and short-range dependencies.
- Do NOT forget about long-range dependencies.
- Try to be as comprehensive as possible.
- Make sure there is a path from earlier sentences to the final answer.
Output Format:
Return a dictionary with one entry per sentence, where each entry has:
- the sentence index (as the key, converted to a string),
- a dictionary with:
- "function_tags": list of tag strings
- "depends_on": list of sentence indices, converted to strings
Here is the expected format:
"1": {
"function_tags": ["problem_setup"],
"depends_on": [""]
"4": {
"function_tags": ["plan_generation"],
"depends_on": ["3"]
"5": {
"function_tags": ["fact_retrieval"],
"depends_on": []
"9": {
"function_tags": ["active_computation"],
"depends_on": ["4", "5"]
"24": {
"function_tags": ["uncertainty_management"],
"depends_on": ["23"]
"32": {
"function_tags": ["final_answer_emission"],
"depends_on": ["9, "30", "32"]
Here is the math problem:
<PROBLEM>
Here is the full chain-of-thought, broken into sentences:
<SENTENCES>
Now label each sentence with function tags and dependencies.
Appendix C Sentence category probing
We trained a linear classifier to identify sentence categories based on activations. We employed a multinomial logistic regression with L2 regularization ($C=1.0$) on the residual stream activity from layer $47$ (last layer) of R1-Distill-Qwen-14B. For evaluating accuracy, we implemented a group-5-fold cross-validation that ensured examples from the same problem response remained in either the training or testing set to prevent data leakage. We averaged the residual stream activity across tokens to create sentence-level representations, whose dimensions were then standardized. To address class imbalance in the training data, we employed balanced class weights. The model demonstrated strong discriminative power across all reasoning categories, achieving a macro-F1 score of $0.71$. The confusion matrix presented in Figure 8 reveals high classification accuracies for categories such as active computation ($0.74$), uncertainty management ($0.79$), and problem setup ($0.83$), while showing some confusion between functionally related categories.
Figure 8: Confusion matrix showing the sentence category classification performance of a logistic regression probe trained on activations from layer 47 of the R1-Distill-Qwen-14B model. Values represent the proportion of examples from each true category (rows) classified as each predicted category (columns). Diagonal elements indicate correct classifications.
Appendix D Other reasoning model
D.1 Measuring counterfactual influence
To assess the generalizability of our counterfactual importance findings, we replicated our resampling methodology on R1-Distill-Llama-8B, applying the same experimental parameters (e.g., temperature = $0.6$ and top-p = $0.95$) used for R1-Distill-Qwen-14B. We collected 100 rollouts for 10 correct and 10 incorrect base solutions using the identical question set described in Section 2.
The resampling accuracy trajectories for R1-Distill-Llama-8B (Figure 9) demonstrate patterns that are similar to those observed in R1-Distill-Qwen-14B (Figure 2). Specifically, we observe similar characteristic accuracy fluctuations throughout the reasoning traces, with notable spikes and dips occurring at sentences corresponding to critical reasoning transitions.
Figure 10 shows that R1-Distill-Llama-8B exhibits similar sentence category effects whereby plan generation and uncertainty management sentences demonstrate higher counterfactual importance compared to active computation and fact retrieval sentences (see Figure  3 for R1-Distill-Qwen-14B).
This cross-model validation supports our claim that reasoning traces are structured around high-level organizational sentences rather than low-level computational steps. The consistency of counterfactual importance patterns suggests that our sentence-level attribution framework captures fundamental properties of chain-of-thought reasoning that generalize beyond specific model implementations.
Figure 9: Accuracy over 100 rollouts at each sentence for (A) one correct and (B) one incorrect base solution for R1-Distill-Llama-8B. Red dots mark significant spikes or dips. Local minima and maxima sentences are annotated with category initials. Our analyses focus on the counterfactual KL-divergence between sentences, but resampling accuracy is visualized here as it is more intuitive.
Figure 10: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance for R1-Distill-Llama-8B, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown.
D.2 Attention aggregation
R1-Distill-Llama-8B displayed receiver-head patterns largely consistent with those of R1-Distill-Qwen-14B. The histogram of attention heads’ vertical-attention scores displays a right tail, indicating that some attention heads tend to particularly focus attention on a subset of sentences (Figure 12A). Interestingly, the R1-Distill-Qwen-14B receiver-heads tended to be more frequent in later layers (see below, Figure 17), which was not evident in R1-Distill-Llama-8B (Figure 11).
The R1-Distill-Qwen-14B and R1-Distill-Llama-8B receiver heads displayed consistent patterns related to sentence types, such that plan generation, uncertainty management, and self checking sentences received heightened attention; although visually, the differences to fact retrieval and active computation may be less prominent, paired t-tests (paired with respect to a given response) showed that plan generation, uncertainty management, and self checking always significantly surpassed fact retrieval and active computation (six paired t-tests: $p$s .$\leq.01$).
No R1-Distill-Llama-8B results are provided for the attention suppression analysis, as that method was principally used for the case study, and no new case study was performed for R1-Distill-Llama-8B.
Figure 11: The plots here show the vertical-attention score patterns associated with the R1-Distill-Llama-8B data. A. This histogram shows the kurtosis values across all attention heads, averaged across all reasoning traces; parallels Figure 4 based on the R1-Qwen-14B data. B. This scatterplot shows the kurtosis of each head’s vertical-attention score, organized by layer. Figure 17 is the R1-Distill-Qwen-14B version of this figure, which showed an upward trend into later layers that is not evident here.
Figure 12: Based on the R1-Distill-Llama-8B data, the boxplot shows the average top-64 receiver-head score for each sentence type. The boxes correspond to the interquartile range across different reasoning traces. Figure 5 is the R1-Distill-Qwen-14B version of this figure; note that for the R1-Distill-Qwen-14B figure, the top-16 heads were used. We found that for Llama 8B, examining the top-64 heads yielded more pronounced differences, although the sentence types with the highest scores remain the same.
Appendix E Embeddings model
We used all-MiniLM-L6-v2 with a maximum sequence length of $256$ tokens and a hidden dimension of $384$ as our sentence embeddings model from the sentence-transformers [27] library. We picked a cosine similarity threshold of $0.8$, which is the median similarity value between all sentence removed (i.e., original sentence) and sentence resampled pairs in our dataset.
Appendix F Additional resampling results
Figure 13 presents mean counterfactual importances across all eight taxonomic categories for R1-Distill-Qwen-14B, extending the main text results (Figure 3) which showed only the five most frequent sentence types. The expanded view includes three additional categories with lower frequencies. Problem setup sentences occur predominantly at trace beginnings (mean normalized position $\approx 0.1$) with moderate-high counterfactual importance. Self checking sentences tend to occur in the second-half of the traces and show lower counterfactual importance. Final answer emission sentences appear late in traces (mean normalized position $\approx 0.9$) and show the lowest counterfactual importance. The patterns observed in the five-category analysis remain consistent when examining the full taxonomy.
Figure 13: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance for R1-Distill-Qwen-14B, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. All sentence types are shown.
Appendix G Additional resampling details
G.1 Counterfactual versus resampling importance
The resampling importance metric introduced in Section 3.2 treats all resampled sentences as equally informative, but different sentence types may exhibit varying degrees of overdetermination during resampling. Overdetermination occurs when resampled sentences $T_{i}$ are frequently similar to the original sentence $S_{i}$ (i.e., $T_{i}\approx S_{i}$), indicating that the reasoning context strongly constrains what can be expressed at that position. We present empirical evidence that counterfactual importance is a more nuanced measure by accounting for semantic divergence in resampled content.
Some sentences are more overdetermined than others. Figure 14A shows that uncertainty management and plan generation sentences produce semantically different alternatives in a large proportion of resamples, while active computation and problem setup sentences show lower divergence rates.
The transition matrix in Figure 14B shows how sentence categories change under resampling. For instance, uncertainty management and active computation sentences are usually replaced by sentences of the same category, whereas plan generation and fact retrieval sentences are more often resampled into a variety of other categories.
Figure 14: (A) Fraction of semantically different resampled sentences by category, showing that uncertainty management and plan generation sentences produce more divergent alternatives when resampled. (B) Transition probabilities between original and resampled sentence categories.
These resampling behaviors create systematic differences between our counterfactual and resampling importance metrics. Figure 15 demonstrates that the relationship between the two metrics varies substantially across sentences and sentence categories. The counterfactual importance metric aims to address overdetermination by explicitly filtering for semantically different resamples, providing a more targeted measure of causal influence. In contrast, the resampling metric potentially overestimates the importance of sentences that consistently produce similar content when resampled.
Figure 15: Comparison between counterfactual and resampling importance metrics across sentence categories. Each point represents a single sentence and the dashed gray line is the $y=x$ line.
However, the counterfactual importance metric can yield high-variance estimates when the number of semantically divergent resampled sentences is low (e.g., $<10$), as the conditional probability estimates become less reliable with limited data. Alongside the limitations discussed in Section 8, this represents another constraint of our approach that future work should investigate further.
G.2 Sentence-to-sentence importance
Beyond measuring individual sentence importance, our framework quantifies causal dependencies between specific sentence pairs within reasoning traces. Figure 16 displays the sentence-to-sentence importance matrix for problem #2236 (incorrect) (“Each page number of a 488-page book is printed one time in the book. The first page is page 1 and the last page is page 488. When printing all of the page numbers, how many more 4’s are printed than 8’s?”), showing how individual sentences influence downstream reasoning steps. Below we list a few illustrative cases.
  • 12-PG $\rightarrow$ 16-PG. The planning in sentence 12 (“1. Count the number of 4’s in the units place across all page numbers”) raises the probability that the model produces sentence 16 (“Starting with the 4’s.”) by $0.39$. A plan statement triggers a subordinate planning step.
  • 8-FR, 9-PG, 12-PG, 14-PG $\rightarrow$ 32-UM. The uncertainty management in sentence 32 (“However, I need to check if 440-449 is fully included.”) receives sizeable positive influence from several earlier sentences: 8-FR ($+0.11$), 9-PG ($+0.06$), 12-PG ($+0.12$), 14-PG ($+0.10$). This forms the dense horizontal band at row index 32.
  • 39-RC ↛$\not\rightarrow$ 83-UM. The result consolidation in sentence 32 (“Now, summing up all the 4’s: - Units: 48 - Tens: 50 - Hundreds: 89. Total 4’s = 48 + 50 + 89 = 187.”) decreases the likelihood (i.e., inhibits) of 83-UM (“Wait, but just to be thorough, let me check the hundreds place for 4’s again.”) by $0.22$.
  • 52-AC ↛$\not\rightarrow$ 65-SC. The computation in sentence 52 (“The first four blocks 80-89, 180-189, 280-289, 380-389 each contribute 10 eights in the tens place.”) decreases the likelihood of 65-SC (“Let me go through each step again to make sure I didn’t make a mistake.”) by $0.16$.
  • 63-AC $\rightarrow$ 64-UM, 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The computation in sentence 63 (“So, the difference is 187 – 98 = 89.”) propagates forward, increasing the likelihood of 64-UM ($+0.24$), 65-SC ($+0.17$), 69-SC ($+0.16$), 75-SC ($+0.28$), 83-UM ($0.23$), and 86-SC ($0.16$). This forms the dense vertical band originating from column index 63.
  • 64-UM $\rightarrow$ 65-SC, 69-SC, 75-SC, 83-UM, 86-SC. The uncertainty management in sentence 64 (“Wait, that seems quite a large difference.”) further amplifies the same downstream block: 65-SC ($+0.32$), 69-SC ($+0.25$), 75-SC ($+0.26$), 83-UM ($0.25$), and 86-SC ($0.25$).
  • 83-UM $\rightarrow$ 86-SC, 90-FAE. Even very late checks matter. Sentence 83 (“Wait, but just to be thorough, let me check the hundreds place for 4’s again.”) increases the chance of 86-SC (“Correct. And for the tens place…”) by $0.43$ and of the final answer in 90-FAE by $0.41$.
Figure 16: Sentence-to-sentence importance matrix for the 32 most important sentences in problem #2236 (incorrect), selected based on total outgoing and incoming importance. Each cell $(i,j)$ shows the causal importance of sentence $i$ on sentence $j$, calculated as the difference in the probability sentence $j$ semantically occurs ($>0.8$ cosine similarity) when sentence $i$ is present versus resampled.
We provide the following pseudocode for estimating sentence-to-sentence importance:
Input: Sentence index i, target sentence index j (where j > i), threshold t = 0.8
Output: Importance score importance(i -> j)
1. Get rollouts R_keep where sentence i was kept (resampling from i+1)
2. Get rollouts R_remove where sentence i was removed (resampling from i)
3. For each rollout r in R_keep:
   a. Extract all sentences S_r from rollout r
   b. Find best matching sentence to target sentence j:
      - Compute sentence embeddings
      - Calculate cosine similarity between each s in S_r and target j
      - Select sentence with highest similarity if similarity >= t
   c. Add to matches_keep if valid match found
4. For each rollout r in R_remove:
   a. Extract all sentences S_r from rollout r
   b. Find best matching sentence to target sentence j (same process as step 3b)
   c. Add to matches_remove if valid match found
5. Calculate match rates:
   match_rate_keep = |matches_keep| / |R_keep|
   match_rate_remove = |matches_remove| / |R_remove|
6. Return importance(i -> j) = match_rate_keep - match_rate_remove
Appendix H Additional receiver head information
Receiver heads – heads receiving high kurtosis scores – are more common in late layers (Figure 17). Examples of receiver heads are shown in Figure 18, showing how the highest kurtosis head consistently narrows attention on particular sentences, and Figure 19, showing how there exist many heads that narrow attention on particular sentences.
Figure 17: This scatterplot shows the kurtosis of each head’s vertical-attention score, organized by layer. There is an upward trend across layers and a strong uptick among some late-layer heads.
Figure 18: The attention weight matrices for the receiver head with the highest kurtosis score are shown here for all twenty responses. The coloring was defined such that the darkest navy corresponds to values surpassing 99.5th percentile value of each matrix. White is zero.
Figure 19: The attention weight matrices for response #1591 (incorrect) are shown here for the 20 attention heads yielding the highest kurtosis score across all responses. No effort was taken to “cherry-pick” responses showing prominent receiver head patterns; we are showing #1591 (incorrect) because it corresponded to the alphabetically earliest problem number among the ten problems analyzed (correct/incorrect chosen randomly). The coloring was defined such that the darkest navy corresponds to values surpassing 99.5th percentile value of each matrix. White is zero.
Appendix I Reasoning versus base model differences in receiver heads
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models. We submitted the reasoning traces to a base model version of Qwen-14B and identified receiver heads. For both models, we sorted all sentences by their mean receiver-head score using the 16 attention heads with the highest kurtoses. The highest percentile sentences received greater attention by the reasoning model - e.g., the highest-percentile sentences receive 1.8x more attention via top-16 heads in the reasoning model compared to the base model (Figure 20). Additionally, lower percentile sentences receive less attention through the top-16 heads. This conclusion is somewhat tenuous, as no base-model difference is seen when this result is tested using R1-Distill-Llama-8B. Nonetheless, based on the Qwen-14B data, it appears the model has learned to narrow its attention toward particular sentences.
Figure 20: The navy and red lines on the left show the receiver-head scores assigned to sentences, averaged across the 16 heads with the highest kurtoses. The green lines on the right represent the ratio of the navy and blue lines for a given sentence rank. Sentences with high receiver head scores receive more attention in the reasoning model compared to the base model.
Appendix J Effects of ablating receiver heads
To test the causal hypothesis that the receiver heads identified in Section 4 are functionally important for reasoning, we performed an experiment ablating receiver heads and evaluating how this impact’s model accuracy. This intervention is designed to measure the direct impact of removing these heads on task performance and to evaluate the possibility that they may be more important than typical heads.
J.1 Methodology
We continue to use problems from the MATH dataset. We selected 32 problems where the non-ablated model achieves 10-90% accuracy on average. For each problem, we ran R1-Distil-Qwen-14B sixteen times, while allowing the model to output up to $2^{16}$ (16,384) tokens. Responses that did not produce an answer by that point were marked as incorrect.
We compared the effect of ablating 128 attention heads (approx. 7% of all heads), 256 heads (approx. 13%), or 512 heads (approx. 27%). The ablation strategies were:
  1. 1.
    Receiver head ablation: We ablated the top-N heads with the highest average kurtosis scores.
  2. 2.
    Random non-receiver (control) ablation: For each layer where $k$ receiver heads were ablated, we ablated $k$ heads chosen randomly from the set of heads not selected from that same layer. This ensures a matched comparison with no overlap.
Note that receiver heads are more common in late layers (see above, Figure 17). By ensuring that both conditions included an equal number of heads from each layer (rather than selecting 128, 256, or 512 heads randomly across all layers), this ensures that differences cannot be explained simply by differences in the layers selected.
In the 512-head ablation condition, a majority of attention heads in some late layers were marked as receiver heads. For these layers, the non-receiver control condition was modified to ablate the corresponding number of heads with the lowest kurtosis scores to ensure a valid comparison set. For instance, if 60% of layer 43 heads are in the top-512, then the control condition included the 60% with the lowest kurtosis score, meaning that there is 20% overlap for that layer.
J.2 Results and Discussion
Our experiments show that a large number of heads must be ablated to induce a significant drop in performance compared to the baseline level of accuracy (baseline = 64.1%, 95% CI: [56.0%, 72.1%]). Regardless of whether receiver heads or non-receiver heads are targeted, ablating 128 heads produces differences in accuracy that insignificantly differ from baseline accuracy, and ablating 256 heads still produces only a small drop in accuracy (Table 2). [^2] We are not aware of prior studies on attention head ablation for models generating long chain-of-thought reasoning, making it difficult to establish what is a typical number of heads to ablate. Potentially, a large number is necessary because the long reasoning traces (sometimes exceeding 10,000 tokens) provide extensive opportunities for error correction and compensatory computation.
Table 2: Answer accuracy on MATH problems for different self-attention-head ablation conditions. The brackets show the 95% confidence interval for each accuracy estimate.
Heads Ablated Receiver heads Random heads
256 48.8% [39.3%, 58.3%] 52.7% [43.0%, 62.5%]
512 27.7% [17.2%, 38.2%] 37.3% [27.5%, 47.1%]
[…]
Appendix K Receiver head correlations with sentence-sentence resampling importance
A sentence’s receiver-head score predicts its impact on downstream sentences per the resampling method. Specifically, our resampling approach (Section 3) produces a matrix describing each sentence’s counterfactual importance on each future sentence. Based on said matrix, we can compute a sentence-level measure analogous to the receiver head score by taking the average value below the matrix diagonal; as with the receiver head score, we omit the four entries immediately below the diagonal to avoid simply capturing proximal effects and focus on farther downstream impacts; we take the absolute value of the resampling matrix to produce an estimate of how a given sentence may both upregulate or downregulate future sentences.
[…]
Appendix L Case study transcript
Below is a transcript of the model’s response used for the case study (Section 6). The first 74 sentences following <think> are shown. After this point, the model converged on 100% resampling accuracy for all subsequent sentences.
  1. 0.
    (problem_setup): Okay, so I have this problem where I need to find out how many bits the base-16 number 6666616 has when it’s converted to base-2.
  2. 1.
    (uncertainty_management): Hmm, let’s see.
  3. 2.
    (fact_retrieval): I remember that each hexadecimal digit corresponds to exactly 4 binary digits, or bits.
  4. 3.
    (plan_generation): So, maybe I can just figure out how many hexadecimal digits there are and multiply by 4 to get the number of bits.
  5. 4.
    (uncertainty_management): Let me check that.
  6. 5.
    (fact_retrieval): First, the number given is 66666 in base 16.
  7. 6.
    (active_computation): Let me count the digits: 6, 6, 6, 6, 6.
  8. 7.
    (result_consolidation): That’s five digits in total.
  9. 8.
    (active_computation): So, if each digit is 4 bits, then 5 digits would be 5×4=20 bits.
  10. 9.
    (uncertainty_management): Wait, is that always the case?
  11. 10.
    (fact_retrieval): I think so because each hex digit maps directly to 4 bits without any overlap or sharing.
  12. 11.
    (result_consolidation): So, 6666616 would be 20 bits in binary.
  13. 12.
    (uncertainty_management): But wait, I should make sure that this is correct by actually converting the number or at least checking if there’s any leading zero that might affect the bit count.
  14. 13.
    (plan_generation): Alternatively, maybe I can calculate the value of 6666616 in decimal and then find out how many bits that number would require.
  15. 14.
    (plan_generation): Let’s try that approach to cross-verify.
  16. 15.
    (fact_retrieval): So, to convert 6666616 to decimal, I can expand it as follows:
  17. 16.
    (fact_retrieval): Each digit represents a power of 16.
  18. 17.
    (fact_retrieval): Starting from the right, the digits are $16^{0}$,$16^{1}$,$16^{2}$,$16^{3}$,$16^{4}$.
  19. 18.
    (fact_retrieval): So, the number is:
  20. 19.
    (active_computation): 6×$16^{4}$+6×$16^{3}$+6×$16^{2}$+6×$16^{1}$+6×$16^{0}$
  21. 20.
    (plan_generation): Let me compute each term:
  22. 21.
    (fact_retrieval): First, $16^{4}$ is 16×16×16×16.
  23. 22.
    (active_computation): $16^{2}$ is 256, so $16^{3}$ is 256×16=4096, and $16^{4}$ is 4096×16=65536.
  24. 23.
    (active_computation): So, 6×65536=393216
  25. 24.
    (active_computation): Next, $16^{3}$ is 4096, so 6×4096=24576
  26. 25.
    (active_computation): Then, $16^{2}$ is 256, so 6×256=1536
  27. 26.
    (active_computation): $16^{1}$ is 16, so 6×16=96
  28. 27.
    (active_computation): And $16^{0}$ is 1, so 6×1=6
  29. 28.
    (result_consolidation): Now, adding all these up:
  30. 29.
    (active_computation): 393216+24576=417,792
  31. 30.
    (active_computation): 417,792+1536=419,328
  32. 31.
    (active_computation): 419,328+96=419,424
  33. 32.
    (active_computation): 419,424+6=419,430
  34. 33.
    (result_consolidation): So, 6666616 is equal to 419,430 in decimal.
  35. 34.
    (fact_retrieval): Now, to find out how many bits this number requires, I can use the formula for the number of bits needed to represent a number N in binary, which is $floor(log_{2}(N))+1$.
  36. 35.
    (plan_generation): So, let’s compute $log_{2}(419,430)$.
  37. 36.
    (uncertainty_management): Hmm, I don’t have a calculator here, but I can estimate it.
  38. 37.
    (fact_retrieval): I know that $2^{18}$ =262,144
  39. 38.
    (fact_retrieval): $2^{19}$ =524,288
  40. 39.
    (result_consolidation): So, 419,430 is between $2^{18}$ and $2^{19}$.
  41. 40.
    (result_consolidation): Therefore, $log_{2}(419,430)$ is between 18 and 19.
  42. 41.
    (active_computation): The floor of that would be 18, so the number of bits is 18+1=19 bits.
  43. 42.
    (uncertainty_management): Wait, but earlier I thought it would be 20 bits because it’s 5 hex digits each being 4 bits.
  44. 43.
    (result_consolidation): Now, according to this, it’s 19 bits.
  45. 44.
    (uncertainty_management): There’s a discrepancy here.
  46. 45.
    (uncertainty_management): I must have made a mistake somewhere.
  47. 46.
    (self_checking): Wait, let me check my calculations again.
  48. 47.
    (uncertainty_management): Maybe I messed up the decimal conversion.
  49. 48.
    (fact_retrieval): Wait, 66666 in hex.
  50. 49.
    (plan_generation): Let me convert it step by step again.
  51. 50.
    (active_computation): 6×$16^{4}$: 6×65536=393216
  52. 51.
    (active_computation): 6×$16^{3}$: 6×4096=24576
  53. 52.
    (active_computation): 6×$16^{2}$: 6×256=1536
  54. 53.
    (active_computation): 6×$16^{1}$: 6×16=96
  55. 54.
    (active_computation): 6×$16^{0}$: 6×1=6
  56. 55.
    (active_computation): Adding them up: 393216+24576=417,792
  57. 56.
    (active_computation): 417,792+1536=419,328
  58. 57.
    (active_computation): 419,328+96=419,424
  59. 58.
    (active_computation): 419,424+6=419,430.
  60. 59.
    (result_consolidation): That seems correct.
  61. 60.
    (result_consolidation): So, 419,430 in decimal.
  62. 61.
    (fact_retrieval): Now, let’s check $2^{18}$ is 262,144, $2^{19}$ is 524,288.
  63. 62.
    (result_consolidation): So, 419,430 is less than 524,288, so $log_{2}(419,430)$ is less than 19.
  64. 63.
    (active_computation): So, $floor(log_{2}(419,430))$ is 18, so number of bits is 19.
  65. 64.
    (uncertainty_management): But wait, when I thought about each hex digit being 4 bits, 5 hex digits would be 20 bits.
  66. 65.
    (uncertainty_management): So, why is there a difference?
  67. 66.
    (plan_generation): Ah, perhaps because leading zeros are not counted.
  68. 67.
    (fact_retrieval): So, when converting from hex to binary, each digit is 4 bits, but if the leading hex digits are such that when converted to binary, they don’t add any leading zeros beyond the 4 bits per digit.
  69. 68.
    (uncertainty_management): Wait, but 66666 in hex is 5 digits, so 5×4=20 bits.
  70. 69.
    (result_consolidation): But according to the decimal conversion, it’s 19 bits.
  71. 70.
    (result_consolidation): That suggests that when converted to binary, it’s 19 bits, not 20.
  72. 71.
    (uncertainty_management): So, which is correct?
  73. 72.
    (plan_generation): Wait, perhaps when converting $66666_{16}$ to binary, it’s exactly 20 bits, but the leading bits are zeros, so when counting the number of bits, we don’t count leading zeros.
  74. 73.
    (fact_retrieval): Wait, but in the problem, it’s just asking how many bits it has when written in base 2.
Sentence 13 (“Alternatively, maybe I can calculate the value of $66666_{16}$ in decimal and then find out how many bits that number would require.”) was found to have the highest counterfactual importance among any sentence in the response (see Figure  2A). Five alternative possible sentences resampled at the sentence 13 position, marked by whether they eventually led to a correct or incorrect answer, are as follows:
  1. 1.
    ($\times$) Let me think. The first digit is 6, which in binary is 0110.
  2. 2.
    ($\checkmark$) Let me try converting the number to decimal first…
  3. 3.
    ($\times$) Let me try converting the first few digits to binary to see how it goes.
  4. 4.
    ($\times$) Let me think about the conversion process.
  5. 5.
    ($\checkmark$) Let me try converting the number step by step.
Appendix M Sentence position effects on receiver-head scores
A sentence’s position within the reasoning trace will tend to influence its measured receiver score.
As a reasoning trace progresses, the number of possible broadcasted sentences will necessarily increase. For instance, by sentence 20, there might be only two broadcasted sentences (each receiving 50% of attention from sentences 21-29), whereas by sentence 100, there could be ten broadcasted sentences (each receiving 10% of attention from sentences 101-109). As the sum of an attention weight row will sum to 1 (at the token level), later sentences will distribute their attention across a larger number of past sentences. This dilution of attention creates downward pressure on the receiver-head scores of later sentences. This is the case even though a receiver head score extends through all subsequent low-competition or high-competition periods. For example, broadcasting sentence 20 will face limited competition from receiving sentence 21-29 attention and high competition for sentences 101-109, whereas broadcasting sentence 100 will exclusively face high competition, pushing its score downward as broadcasting-sentence position increases.
There also exists a proximity effect on receiver-head scores that operates in the opposite direction of the above effect. Although broadcasted sentences are attended by all subsequent sentences to some degree, this will be more so the case for more recently subsequent sentences (e.g., receiving more attention from a sentence 10 sentences downstream than one 20 sentences downstream). For sentences late in the reasoning trace, the average distance to future sentences will be shorter. For example, if a reasoning trace contains 120 sentences, then sentence 100 will be at most 19 sentences apart from any given future sentence, whereas sentence 20 will be at most 99 sentences apart. To a degree, the analyses in the report account for proximity effects by ignoring the 4 sentences immediately proximal to a given sentence when calculating vertical-attention scores. However, this will not fully address proximity effects.
We see no reason why the downward pressure of sentence position on receiver-head scores (attention dilution) will be equal in magnitude to the upward pressure of sentence position (proximity effects).
For the preparation of the present report, we conducted exploratory analyses evaluating whether the above confounding factors invalidate any presented finding, and we did not find evidence that this is the case. Thus, rather than pursuing some technique to account for the above pressures (e.g., linearly weighing attention weight matrices based on their position), we opted to only account for these factors in a minimal fashion by ignoring the attention among sentences just 4 sentences apart.